### v2

# CP-Agent: Agentic Constraint Programming

Abstract
Translating natural language into formal constraint models requires expertise in the problem domain and modeling frameworks. To investigate whether constraint modeling benefits from agentic workflows, we introduce CP-Agent, a Python coding agent using the ReAct framework with a persistent IPython kernel. Domain knowledge is provided through a project prompt of under 50 lines. The agent iteratively executes code, observes the solver’s feedback, and refines models based on the execution results.
We evaluate CP-Agent on CP-Bench’s 101 constraint programming problems. We clarified the benchmark to address systematic ambiguities in problem specifications and errors in ground-truth models. On the clarified benchmark, CP-Agent solves all 101 problems. Ablation studies indicate that minimal guidance outperforms detailed procedural scaffolding, and that explicit task management tools have mixed effects on focused modeling tasks.
1 Introduction
Constraint Programming (CP) is a declarative paradigm for solving combinatorial problems where users specify what constraints a solution must satisfy, rather than prescribing how to find it [26]. The primary limitation preventing wider adoption of CP remains the modeling bottleneck: translating natural language problem descriptions into formal constraint models requires deep expertise in both the problem domain and the modeling framework [5]. This challenge has been characterized as the “holy grail” of constraint programming [24].
Large Language Models (LLMs) are now being applied to automate this translation process. The CP-Bench benchmark [14], introduced to evaluate LLM-driven constraint modeling capabilities, contains 101 diverse combinatorial problems sourced from various established CP community resources. These problems encompass optimization and satisfaction tasks, representing a variety of distinct constraint types. This benchmark addresses limitations of existing datasets, such as NL4Opt [17] and Logic Grid Puzzles [25], which lack the diversity and complexity necessary for evaluating constraint modeling capabilities.
The CP-Bench study [14] evaluated multiple LLMs across three modeling frameworks. With direct prompting, the best models achieved up to 65% accuracy using Python-based frameworks (CPMpy, OR-Tools), while MiniZinc (a popular constraint modeling language) peaked around 50%. Applying inference-time methods (repeated sampling combined with self-verification) to CPMpy improved accuracy to approximately 70%.
The choice between fixed and agentic workflows is well-established in the literature. As Wei et al. [29] demonstrated with chain-of-thought prompting, and subsequent work by Yao et al. [32] and Schick et al. [19] confirmed that one should use a fixed, single-shot workflow for well-scoped, short-horizon tasks, and prefer an agentic, tool-using loop as task complexity, multi-step reasoning, or external information needs increase. We hypothesize that constraint modeling from natural language falls into the latter category, requiring an agentic workflow with iterative refinement, hypothesis testing, and dynamic adaptation based on solver feedback. Our results confirm this hypothesis.
To test this hypothesis, we introduce CP-Agent, a constraint modeling agent based on our general-purpose agentic-python-coder system. CP-Agent operates within a persistent IPython kernel [16], enabling stateful and iterative development. Rather than embedding domain-specific logic into its architecture, the agent is guided by a concise project prompt. This design allows the agent to execute code, observe runtime feedback from the constraint solver, and adjust models based on execution results.
The underlying coding agent is also available as an MCP server [1]. The Model Context Protocol (MCP), now governed by the Linux Foundation, has become a de facto standard for connecting language models with external tools [8]. This integration enables interactive constraint modeling within MCP-compatible clients such as Claude Desktop or Cursor, where users can guide the agent through complex modeling decisions.
On the CP-Bench dataset of 101 problems (with added clarifications for some ambiguous problems and output format requirements), CP-Agent solves all problems correctly. While our primary reference language model is Claude Sonnet 4.5, we also tested this approach with several other models from four different providers. Swapping only the project prompt enables the agent to adapt to new domains, as confirmed by follow-up work [23] on logic programming modeling tasks.
We provide an experimental analysis of an explicit task management tool within our agentic coding workflow. The results are mixed: while task tracking helped on some previously failing problems, it also broke problems that worked without it. The net effect was negative (6 fewer problems solved) with significantly increased token usage, suggesting that for focused, single-file generation tasks, the overhead can distract as often as it organizes.
2 Related Work
CP-Agent aligns with the emerging neurosymbolic paradigm where neural language understanding is combined with symbolic constraint satisfaction [18]. The LLM-based agent translates natural language into formal constraints, while the CP solver provides a sound and complete search over the solution space.
Recent work has explored LLM agents for software engineering tasks. SWE-agent [31] resolves GitHub issues through a file-system-based Agent-Computer Interface with commands for navigating and editing repository files. CP-Agent focuses instead on single-script generation using a stateful IPython kernel, eliminating file I/O overhead for this class of problems. CodeAct [28] argues for using executable Python code as a unified action space for LLM agents, showing that this approach outperforms more restrictive formats. Our work applies this principle to constraint modeling, demonstrating that a general action space can solve specialized tasks by utilizing focused prompts. The iterative refinement in CP-Agent relates to automated program repair work. ChatRepair [30] introduced conversational repair, where an LLM iteratively refines patches based on test failures. RepairAgent [3] treats repair as an autonomous task with tools for code search, patching, and testing. Our agent performs similar iterative refinement but for generative tasks, translating natural language to constraint programs, rather than repairing existing code.
Benchmark quality affects evaluation validity. EvalPlus [13] demonstrated that popular benchmarks contain insufficient tests, with augmented test suites revealing previously undetected errors. ClassEval [4] addressed structural complexity through manual benchmark curation. Our work on CP-Bench addresses problem specification ambiguity, enabling more rigorous measurement of constraint modeling abilities.
3 Agentic Python Coder
CP-Agent is based on a generic Python coding agent equipped with a special prompt for constraint modeling. The coding agent implements a minimal yet capable coding agent using the ReAct framework [32]. The system architecture separates domain-agnostic agent capabilities from task-specific knowledge, enabling rapid adaptation to new problem domains through prompt engineering rather than code modification. The implementation of the generic coding agent is available from a GitHub repository and via PiPy for installation via “pip install agentic-python-coder” [20].
3.1 Architecture
Virtual execution space
Unlike software engineering agents that navigate complex file trees, this architecture targets single-script development, where all state is contained in a single file. The agent operates in a virtual execution space provided by the IPython kernel [16], which maintains a persistent state across multiple code executions. Variables, functions, and imports remain accessible throughout a session, eliminating file I/O overhead while enabling rapid hypothesis testing. The agent operates interactively via a Read–Eval–Print-Loop (REPL), and benefits from immediate feedback to all code manipulations. This design is well-suited for constraint programming models and similar focused generation tasks.
To manage this persistent state and guide the execution flow, we employ LangGraph [12]. This framework manages the agent’s workflow, enabling the development of stateful agents through durable execution and memory management. The framework supports the ReAct pattern [32] of interleaved reasoning and acting, allowing the agent to interact with the execution environment and adjust solutions through iterative cycles. The agent consists of three main components: the ReAct loop controller, which operates the reasoning-action cycle; a minimal tool suite focused on code generation and execution; and a project prompt that provides domain-specific guidance. The ReAct loop (Figure 1) manages the iterative cycle of reasoning about the task, selecting and executing tools, and observing results until the agent produces a final solution.
TaskReasonAct (Tools)Solutionexecutecompleteobserve
Figure 1: The ReAct (Reason and Act) framework employed by the agent.
Tool suite
The agent operates with a minimal, focused tool set. The core tool is python_exec(code: str), which executes Python code in the persistent kernel environment and returns stdout/stderr for solver feedback. Each execution builds on the previous state, allowing the agent to develop solutions incrementally. Calling save_code signals completion and writes the final solution to disk; a maximum iteration limit prevents runaway execution.
Kernel implementation
The system launches an IPython kernel as a separate subprocess, providing process isolation while enabling state persistence. All Python state resides in this long-running kernel, creating a stateful coding environment similar to an interactive session where each execution builds upon previous ones.
Dynamic dependency injection
Rather than having the agent execute pip install statements for adding required Python libraries or relying on pre-built container images, the system implements a custom kernel manager that wraps the standard IPython kernel launch. Using the uv package manager [2], the system injects requested packages directly into an ephemeral environment at kernel startup. This approach provides process-level isolation: each session runs in a clean environment with only the specified dependencies, while the kernel maintains state persistence across executions. The design decouples the agent’s reasoning from dependency mechanics, thereby eliminating the latency associated with runtime package installation.
Task management
The agent can use a task management tool to create and maintain an in-memory list of tasks, commonly referred to as a “todo list.” The tool call uses the format todo_write(todos: List[Dict[str, Any]]). Each task must contain id, content, status, and priority fields, with only one task marked as “in_progress” at any time.
LLM integration
The agent connects to language models through OpenRouter, [^1] https://openrouter.ai providing access without direct API dependencies. This approach allows switching between different models by changing a single configuration parameter. The system streams responses with usage statistics and detailed tracking of tool calls. These logs provide insights into the agent’s problem-solving process and resource consumption.
Prompt architecture
The agent operates with a three-layer prompt hierarchy. The system prompt (approximately 100 lines) defines general coding agent behavior and tool usage patterns without task-specific specialization, instructing the agent on how to reason about code, execute iterative development cycles, and utilize available tools effectively. The project prompt provides domain-specific knowledge (such as constraint modeling), examples, and best practices relevant to a particular problem class. The task prompt contains the specific problem to solve, typically read from a task.md file in the working directory or provided directly via a command line parameter. This separation enables the same agent infrastructure to tackle diverse problem domains by swapping project prompts while maintaining consistent operational behavior.
4 CP-Agent
For constraint programming, the generic coding agent is equipped with a domain-specific project prompt and uses CPMpy (see below) as the modeling framework. The approach relies on the agent’s ability to use general coding capabilities guided by domain expertise expressed in the project prompt, rather than specialized constraint programming logic embedded in the architecture. The agent can test simplified problem instances, verify constraint satisfaction on example data, implement custom validation logic, and refine models based on solver feedback. This flexibility, combined with the persistent execution environment, enables the agent to discover and correct modeling errors that would escape detection in single-pass generation approaches.
4.1 CPMpy
The agent uses CPMpy [7] as the constraint programming foundation. This Python-based library builds on NumPy for efficient handling of decision variables as arrays, allowing integration with the broader Python ecosystem. Its solver-agnostic design enables modeling problems once and solving them using various backends, including CP, MIP, SMT, and SAT solvers. The choice of Python matches findings from Michailidis et al. [14] that LLMs perform better with Python-based frameworks than domain-specific languages, likely due to the billions of lines of Python code in their training data.
4.2 Project Prompt
The project prompt (cpmpy.md) is under 50 lines and provides focused guidance for constraint modeling. The complete prompt is included in the GitHub repository [20].
The prompt contains five core rules: use CPMpy’s declarative modeling rather than implementing search algorithms, output only valid JSON using json.dumps(), always import the json module when outputting JSON, check the exact output format required, and manually test solutions to ensure they satisfy the problem requirements. A basic code template demonstrates the standard structure, and brief sections cover three essential CPMpy constraints (AllDifferent, sum, Circuit) and optimization via model.minimize() and model.maximize().
This minimal prompt resulted from an ablation study. Initial experiments utilized a substantially longer prompt, approximately 800 lines in length, which contained detailed modeling procedures, extensive problem-type catalogs, and debugging guidance. The shorter version achieved equal or better performance, suggesting that for modern LLMs with constraint programming knowledge in their training data, concise guidance focusing on common pitfalls suffices—excessive detail may limit problem-solving flexibility.
5 Benchmark Problems
The CP-Bench dataset [14] provides a benchmark for evaluating automated constraint modeling capabilities. It contains 101 constraint programming problems drawn from four established sources: CSPLib [6], CPMpy examples [7], models by Kjellerstrand [11], and problems from the APLAI course at KU Leuven. The problems span diverse combinatorial classes, including scheduling, assignment, sequencing, graph problems, and combinatorial design, with 30 optimization problems and 71 decision problems providing balanced coverage of constraint programming challenges.
Each problem is presented as a natural language description in Markdown format, along with input data and required output variables. Solutions must execute within a specified time limit and produce results in JSON format, using the specified variable names. CP-Bench includes ground-truth CPMpy models for all problems, enabling semantic validation through constraint satisfaction checking rather than relying solely on syntactic correctness or execution success. This semantic validation is essential for rigorous evaluation of automated modeling approaches. The ground-truth models are withheld from the agent and used solely for post-hoc verification of generated solutions.
5.1 Fixed Workflow Results
Michailidis et al. [14] evaluated fixed workflow approaches on CP-Bench, testing progressively detailed prompts (basic instructions, modeling guidelines, API documentation) and retrieval-augmented in-context learning. The highest accuracy of 70% was achieved by combining repeated sampling with self-verification using GPT-4-turbo. These fixed workflow approaches share a common limitation: they cannot test hypotheses, debug errors, or verify partial solutions.
5.2 Benchmark Corrections and Clarifications
Our initial evaluation attempts revealed issues with the original benchmark that prevented rigorous automated evaluation. The CP-Bench authors independently acknowledged222personal communication, August 2025 several problems, including ground-truth models with symmetry-breaking constraints that reject valid solutions, unclear output format requirements, and ambiguities in problem statements. Without clear specifications, evaluation conflates the ability to model constraint satisfaction problems with the ability to guess intended formats and interpret ambiguous descriptions. We added clarifications to isolate the actual modeling capability. These clarifications are necessary to maintain focus on constraint modeling rather than resolving ambiguity or making guesses.
Our benchmark updates fall into three categories. First, we clarified 31 ambiguous problem statements. For instance, Problem 080 (TSP) required specifying whether Euclidean distances should be rounded or truncated—different methods lead to different distance matrices and thus different optimal solutions. For Problem 013 (magic square), we corrected a misleading description that stated the goal is to find a “magic square” when the problem actually requires a “magic sequence.” Problem 060 (crossword) provided a visual grid and word list, but required explicit specification of slot lengths and intersection constraints to be solvable. Second, we added structured JSON schemas to all problems to clarify the required output format. Third, we corrected errors in 19 ground-truth models: removing symmetry-breaking constraints from 10 and fixing logic errors in 9 others.
The modifications do not make the constraint modeling problems easier—they remove confounding factors that obscure actual modeling capability. The underlying combinatorial structure and search space complexity remained unchanged; no changes hint at constraint formulations or modeling strategies. The success rates reported in the original CP-Bench paper were obtained using the original benchmark and ground-truth models. Direct comparison between our results and theirs would be inappropriate, as they measure different capabilities. With clear specifications, other methods would likely show higher success rates as well. We document all changes in supplementary materials and provide the clarified benchmark for fair comparison in future work.
6 Experiment
6.1 Experimental Setup
Our experimental setup executed all 101 problems using CP-Agent, implemented as agentic-python-coder (version 2.2.1), with Claude Sonnet 4.5 (version 20250929). Each problem execution produces a solution file containing the generated CPMpy model and a conversation log recording the complete agent interaction. We validated the solutions by running them against the ground-truth CPMpy models. For satisfaction problems, we verified that the solutions satisfy all ground-truth constraints. For optimization problems, we also compared the objective value to confirm optimality. This approach correctly handles problems with multiple valid solutions while ensuring correctness and accuracy. We ran each experiment three times to account for stochastic variance.
The agentic-python-coder and all experimental data, including problem specifications, ground-truth models, generated solutions, and execution logs, are available online [22]. The data is encrypted (password Ckc1ADsKXRfmzD2) to prevent training data contamination.
6.2 Results
CP-Agent achieved 100% accuracy on all 101 CP-Bench problems, successfully solving all problems across all three independent runs. All generated solutions were validated against the reference CPMpy models and produced correct and optimal solutions.
Table 1: Tool usage, token consumption, and execution time for all 101 CP-Bench problems (averaged over three independent runs). Column ex shows the number of python_exec calls (iterative refinement cycles). Columns in/out show input/output tokens in thousands. Column time shows execution time in seconds (rounded).
Problem ex in out time Problem ex in out time
001_car_seq 2.7 24 2 43 052_circular_t 3.7 29 2 47
002_template 2 21 2 52 053_clock_trip 1 14 1 26
003_autocorr 3.3 26 2 44 054_cmo_2012 3 25 3 49
004_golomb 2.3 22 2 42 055_coin3_appl 2 18 2 34
005_all_inter 1.7 19 2 32 056_coins_grid 4.3 38 2 47
006_vessel 2.3 26 3 51 057_contractin 1 14 1 25
007_perfect_sq 4 30 2 44 058_covering_o 2 24 2 42
008_social_gol 7.7 54 3 61 059_crew 2.3 36 4 62
009_basketball 23.7 256 12 245 060_crossword 4 40 3 55
010_nonogram 8 90 6 110 061_crypta 4.7 67 4 74
011_progressive 8 79 5 93 062_eighteen_h 1 12 1 21
012_schurs 2.7 23 2 36 063_facility_l 3 29 3 57
013_magic_sq 1 15 1 22 064_fifty_puzz 1.3 15 1 23
014_crossfig 28.3 737 33 542 065_three_coin 1 15 1 27
015_langford 1.7 19 2 36 066_three_sum 1 13 1 20
016_sports 8.3 75 5 99 067_twelve_pac 1.3 14 1 25
017_bibd 2.7 24 2 37 068_bus_schedu 1 14 1 27
018_word_des 6.3 47 3 80 069_jobshop 3.3 29 3 52
019_steiner 3.7 29 2 41 070_knapsack 1 13 1 22
020_num_part 2.7 21 2 60 071_mario 7.3 86 6 110
021_diamond 4.3 41 4 77 072_minesweepe 5 37 3 57
022_graceful 1 16 1 26 073_n_puzzle 10.7 106 6 119
023_n_queens 2 20 2 42 074_packing_re 8.3 75 5 83
024_costas 2 20 1 30 075_resource_c 5.7 68 4 78
025_hadamard 4.3 31 2 46 076_room_assig 2 21 2 38
026_abbots_puz 1 13 1 23 077_send_more 1 14 1 25
027_added_corn 2 20 2 37 078_set_game 3.3 34 3 60
028_ages_of_th 4 27 2 53 079_sudoku 2 23 2 39
029_age_changi 4.7 38 3 57 080_tsp 3.3 24 2 40
030_allergy 2 21 2 44 081_who_killed 2 21 2 42
031_among 2 17 1 42 082_wolf_goat 3.3 28 3 58
032_appointmen 1 14 1 36 083_zebra 3.7 38 4 59
033_archery_pu 2.3 18 1 49 084_bank_card 1 13 1 23
034_arch_frien 2 24 3 56 085_climbing_s 1.3 16 1 31
035_assignment 2.7 22 2 49 086_color_simp 2 18 1 30
036_autoref 4.3 31 2 46 087_exodus 5 45 4 71
037_bales_of_h 4 33 3 66 088_farmer_and 2.3 19 2 34
038_bananas 1.7 18 2 39 089_five_floor 1 14 1 24
039_best_host 1.7 18 2 33 090_grocery 4.3 31 2 51
040_big_bang2 1.7 21 2 39 091_guards_and 2 20 2 37
041_bin_packin 1 13 1 23 092_hardy_1729 1.7 15 1 28
042_birthday_c 1 14 1 26 093_kidney_exc 3.3 27 2 44
043_bowls_and 1.7 16 1 30 094_magic_sq 1.3 15 1 28
044_broken_wei 2 19 2 40 095_maximal_in 1.7 19 2 36
045_building_b 2.3 20 2 34 096_money_chan 1 13 1 24
046_cabling 1.3 17 2 36 097_movie_sche 2.7 23 2 41
047_calvin_puz 2.7 24 2 49 098_people_in 1 15 2 36
048_candies 1.3 16 2 28 099_subsets_10 1 14 1 27
049_capital_bu 1 14 1 25 100_subset_sum 2.7 21 1 37
050_chess_set 6.3 79 3 59 101_thick_as_t 2.7 26 3 54
051_circling_s 2.3 25 3 47
Token consumption averaged 37k input and 3k output tokens per problem. Problem 014 (crossfig) represents an extreme outlier, consuming 737k input tokens due to its complex crossword-figure structure, which requires extensive iterative constraint refinement.
Execution times ranged from 20 seconds to 542 seconds, with an average of 52 seconds per problem across the three runs. The timing variation between the three independent runs (averaging 48, 57, and 52 seconds, respectively) demonstrates the stochastic nature of LLM-based systems.
Table 1 summarizes computational cost; solution quality is perfect across all runs. The agent called python_exec between 1 and 28 times per problem, with simpler problems requiring fewer iterations and complex scheduling or optimization problems requiring more exploration. Problem 014 (crossfig) required the most Python executions (28 average), followed by 009 (basketball) with 24 executions, reflecting their combinatorial complexity.
To illustrate how the agent approaches constraint modeling, we examine two representative runs that demonstrate different problem-solving patterns.
6.3 Problem 061: Cryptarithmetic Puzzle
The cryptarithmetic problem required assigning unique digits (0–9) to letters in a 20-digit addition equation, ensuring the arithmetic holds while each letter maps to a distinct digit. The agent’s solution process demonstrates strategic adaptation when an initial approach proves unworkable.
The first four executions built the foundation: defining ten integer variables (one per letter), adding an AllDifferent constraint, mapping variables to their positions in the three numbers, and constraining leading digits to be non-zero. This setup proceeded smoothly, establishing the basic problem structure.
The fifth execution attempted to model the problem as a single unit by treating each 20-digit number as a single large integer variable. The agent constructed expressions like num1 = letter[0]*10**19 + letter[1]*10**18 + ... for all three numbers and added a constraint num1 + num2 == result. When the model tried to solve, the solver raised a TypeError: __init__(): incompatible constructor arguments because variable domains with bounds near $10^{20}$ exceeded the solver’s integer representation capacity.
Recognizing the fundamental incompatibility between this modeling approach and solver limitations, the agent abandoned the large-integer strategy entirely. The agent redesigned the model, introducing binary carry variables and modeling addition column by column, thereby mimicking manual arithmetic:
carries = intvar(0, 1, shape=21)
for pos in range(20):
digit1 = letter_to_var[num1_chars[19-pos]]
digit2 = letter_to_var[num2_chars[19-pos]]
result_digit = letter_to_var[result_chars[19-pos]]
carry_in = carries[pos]
carry_out = carries[pos+1]
model += (digit1 + digit2 + carry_in
== result_digit + 10 * carry_out)
This column-wise formulation ensured that all intermediate values remained within standard integer bounds, enabling the solver to determine the correct digit assignment.
6.4 Problem 082: Wolf, Goat, and Cabbage
While the above example demonstrates error recovery, this river-crossing puzzle illustrates the agent’s behavior when proceeding without failures. The agent built the temporal model through four distinct phases.
The first phase established state variables: wolf_pos[t], goat_pos[t], cabbage_pos[t], and boat_pos[t], which track locations (shore 0 or 1) across an 8-step timeline. The second phase fixed boundary conditions—all entities start at shore 0 and end at shore 1. The third phase encoded the movement mechanics: the boat crosses at each step, the farmer transports at most one item per trip, and item movements coincide with boat movements. The fourth phase added safety constraints:
for i in range(steps):
model += ~((goat_pos[i] == cabbage_pos[i])
& (goat_pos[i] != boat_pos[i]))
model += ~((wolf_pos[i] == goat_pos[i])
& (wolf_pos[i] != boat_pos[i]))
These constraints forbid the goat and cabbage (or wolf and goat) from sharing a shore without the farmer present. The systematic layering—state, boundaries, movement, and safety—reflects a deliberate construction process where the agent tests each phase before proceeding to the next. The model solved correctly on the first execution.
These two problems reveal complementary behaviors: complete redesigns when facing fundamental obstacles (Cryptarithmetic) and methodical incremental construction when the path is clear (Wolf, Goat, and Cabbage). The agent adapts its approach rather than following a fixed procedure.
7 Model Generalization
To evaluate whether the approach generalizes beyond Claude Sonnet 4.5, we tested six frontier models on ten representative problems with three runs each. The models span four providers: Anthropic (Sonnet 4.5, Opus 4.5, Haiku 4.5), OpenAI (GPT-5.1-codex-mini), xAI (Grok-code-fast-1), and DeepSeek (v3.2). We excluded Google’s Gemini models due to infrastructure issues during the experimental period. All models used a temperature of 0.0 with a 20-minute timeout per problem.
Table 2 shows the results. All tested models achieved a success rate of 80–100% using this architecture. Sonnet 4.5 maintained perfect accuracy, while Opus 4.5 and GPT-5.1 each reached 97%. Haiku 4.5, Grok, and DeepSeek achieved 80–87% accuracy, with most failures occurring on Problem 009 (basketball), although sporadic failures also occurred on other problems. Basketball requires coordinating nine constraint categories, including mirroring schemes, home/away patterns, rival pairings, and date-specific requirements—the highest constraint density in this subset.
Some runs hit the 20-minute timeout before completion. Review of execution logs confirms these agents were not converging toward correct solutions, suggesting the timeout did not cut short otherwise successful attempts.
These results show the architecture generalizes across model families. Success correlates with model capability rather than provider: the agentic workflow provides value at all capability levels tested. Prompt optimization frameworks, such as DSPy [10], could automatically tune the project prompt for specific models, potentially narrowing the performance gap observed with other models.
Table 2: Multi-model evaluation (3 runs each). Numbers indicate successful runs out of 3. Model success rates: Sonnet 100%, Opus/GPT-5 97%, Haiku 87%, DeepSeek 83%, Grok 80%.
Problem
Sonnet 4.5
Opus 4.5
Haiku 4.5
GPT-5.1
Grok
DeepSeek
001_car_seq 3 3 3 3 3 3
002_template 3 2 3 3 3 2
004_golomb 3 3 3 3 2 3
007_perfect_sq 3 3 3 3 3 3
009_basketball 3 3 1 2 0 0
010_nonogram 3 3 3 3 3 2
013_magic_sq 3 3 3 3 3 3
015_langford 3 3 1 3 2 3
020_num_part 3 3 3 3 3 3
037_bales_of_h 3 3 3 3 2 3
8 Task Management Tool
Explicit task tracking for LLM agents—maintaining a mutable list of pending, in-progress, and completed tasks—was introduced by BabyAGI [15] and formalized in plan-and-solve prompting [27]. The todo_write tool in our coder implements this pattern. Our architecture requires only todo_write, not a corresponding todo_read: without history compaction, all previous tool calls remain in the conversation context, so the agent can always see its earlier task updates. To evaluate whether explicit task management benefits constraint modeling, we conducted an ablation study comparing performance with and without this tool.
We used a smaller model (Claude Haiku 4.5) than our default model (Claude Sonnet 4.5) to reveal potential differences. We selected 40 problems, stratified by difficulty: 20 easy problems (solved in 27 seconds or less by Sonnet 4.5) and 20 hard problems (requiring 58 seconds or more). For each problem, we executed three independent runs with todo_write disabled and three with it enabled. Table 3 presents the results.
Table 3: Task management ablation on 40 problems (Claude Haiku 4.5, 3 runs per condition). Columns: p = success (out of 3), tok = tokens (K, averaged), t = time (s, averaged). Bold highlights problems where success changed.
no todo todo no todo todo
Easy Problem p tok t p tok t Hard Problem p tok t p tok t
013_magic_sq 3 15 26 3 39 35 008_social_gol 3 36 36 3 122 74
022_graceful 3 38 32 3 49 42 009_basketball 3 752 303 0 181 119
026_abbots_puz 3 35 29 3 57 40 010_nonogram 3 47 48 3 112 67
041_bin_packin 3 10 13 3 52 48 011_progressive 1 201 111 2 730 197
042_birthday_c 3 19 20 3 52 40 014_crossfig 0 1986 617 2 1592 492
049_capital_bu 3 10 18 3 37 41 016_sports 0 112 70 1 712 228
053_clock_trip 2 17 18 3 87 54 018_word_des 3 41 40 3 160 90
057_contractin 3 16 18 3 38 35 021_diamond 3 69 63 3 130 77
062_eighteen_h 3 15 16 3 57 49 037_bales_of_h 3 40 40 3 149 81
064_fifty_puzz 3 10 12 3 35 29 050_chess_set 3 10 16 3 38 31
065_three_coin 3 15 16 3 70 60 059_crew 3 80 55 0 137 75
066_three_sum 3 9 13 3 31 26 061_crypta 3 43 31 2 297 127
067_twelve_pac 3 9 12 3 39 33 071_mario 3 243 145 2 291 135
068_bus_schedu 3 18 20 3 50 39 073_n_puzzle 3 203 95 3 273 122
070_knapsack 3 9 11 3 37 32 074_packing_re 3 22 25 3 134 81
077_send_more 3 11 13 3 41 37 075_resource_c 3 55 51 3 86 61
084_bank_card 3 10 16 3 26 24 078_set_game 3 24 24 3 105 70
089_five_floor 3 11 14 3 33 30 082_wolf_goat 3 62 60 0 286 156
096_money_chan 3 13 22 3 32 34 083_zebra 3 30 32 3 150 91
099_subsets_10 3 16 19 3 28 26 087_exodus 3 123 88 3 223 106
Easy (60 runs) 59 60 Hard (60 runs) 52 45
On easy problems, todo_write increased token consumption by 191% (from 15K to 44K average) and execution time by 111% (from 18s to 38s), while improving success by just one run (59/60 to 60/60). On hard problems, todo_write reduced success from 52/60 to 45/60 while still adding overhead: 41% more tokens (209K to 295K) and 27% longer execution (98s to 124s).
Examining individual problems reveals the pattern. The task management tool helped three previously failing hard problems—crossfig, sports, and progressive—by providing structure for extended problem-solving. However, it broke three problems that worked without it: basketball, crew, and wolf_goat all dropped from 3/3 to 0/3. The task-tracking overhead appears to distract the agent from solutions it would otherwise find.
9 Discussion
Our experiments yield several key observations regarding the design and evaluation of agentic systems for constraint modeling.
The ablation study, which shows that a short prompt outperforms an 800-line version, suggests that for foundation models with constraint programming knowledge in their training data, the prompt’s role is to set high-level strategy and details on output format rather than provide exhaustive procedural instructions. Prompt compression research supports this finding: shorter, denser prompts can match or exceed verbose prompts [9]. Excessive detail may limit the agent’s ability to solve problems effectively.
The primary evaluation used Claude Sonnet 4.5. Multi-model testing on a representative 10-problem subset confirms that the architecture generalizes: six models from four providers achieved success rates of 80–100%, with performance correlating with model capability.
This pattern extends to built-in planning mechanisms. The optional todo_write tool, intended to help the agent manage tasks, reduced overall success when tested with a smaller model. The effect was asymmetric: todo_write helped on already-failing problems but broke problems that worked without it. For problems where the entire solution fits in a single script, the overhead of maintaining an explicit task list is more distracting than beneficial. When state is fully captured in code and execution output, the agent’s internal reasoning seems to suffice without externalized planning.
10 MCP Server
The Model Context Protocol (MCP) provides a standardized interface for connecting language models with external tools and data sources [8]. Introduced by Anthropic in late 2024 [1], MCP has become a de facto standard for agent-tool interaction, with adoption by major AI platforms including OpenAI and integration into development environments such as Cursor, Claude Code, VS Code (via Cline), Windsurf, and OpenAI Codex CLI. In December 2025, MCP governance transferred to the Agentic AI Foundation under the Linux Foundation.
While our standalone CP-Agent excels at batch processing, interactive development benefits from human guidance through complex modeling decisions. We expose the same persistent IPython kernel as an MCP server (ipython_mcp) for integration with Model Context Protocol-compatible clients such as Claude Desktop or Cursor. The server provides python_exec for code execution, the same tool central to our standalone agent, along with three tools for controlling and monitoring the coding infrastructure: python_reset to clear state and install packages to the virtual environment, python_status to query session information, and python_interrupt to halt long-running computations. This interactive mode also serves an educational purpose: users can observe how natural language constraints are formalized into CPMpy models, request explanations, and refine models based on solver feedback. Once developed, the generated code can be saved for deployment in other contexts.
Our MCP integration complements the system MCP-Solver [21], which provides specialized tools for constraint modeling (MiniZinc, PySAT, Z3, ASP). In contrast, the ipython_mcp server offers general-purpose Python execution, where domain adaptation occurs through prompts rather than tool specialization.
11 Conclusion
CP-Agent solves all 101 problems in the clarified CP-Bench benchmark, demonstrating that agentic workflows with execution feedback can reliably automate constraint modeling from natural language.
The agent architecture combines the ReAct framework with a persistent IPython kernel. Our ablation study shows that a minimal project prompt is sufficient, and that explicit task management tools generally add overhead for focused modeling tasks. However, they can help maintain focus during longer problem-solving sessions. For foundation models with relevant domain knowledge, concise guidance and stateful execution prove more effective than detailed procedural scaffolding or explicit planning mechanisms.
The architecture’s separation of domain expertise (prompts) from execution infrastructure (the agent) allows adaptation to other scientific computing domains. Follow-up work on ASP-Bench [23] demonstrated excellent performance on 128 answer set programming problems, utilizing the identical architecture with only a domain-specific project prompt, thereby confirming that our approach generalizes across declarative programming paradigms.
Acknowledgments
This research was supported by the Austrian Science Fund (FWF) projects 10.55776/P36420 and 10.55776/COE12.

### v3

@@
-CP-Agent: Agentic Constraint Programming
+Stefan Szeider Algorithms and Complexity Group TU Wien, Vienna, Austria
 Abstract
-Translating natural language into formal constraint models requires expertise in the problem domain and modeling frameworks.
-To investigate whether constraint modeling benefits from agentic workflows, we introduce CP-Agent, a Python coding agent using the ReAct framework with a persistent IPython kernel.
-Domain knowledge is provided through a project prompt of under 50 lines.
-The agent iteratively executes code, observes the solver’s feedback, and refines models based on the execution results.
-We evaluate CP-Agent on CP-Bench’s 101 constraint programming problems.
-We clarified the benchmark to address systematic ambiguities in problem specifications and errors in ground-truth models.
-On the clarified benchmark, CP-Agent solves all 101 problems.
-Ablation studies indicate that minimal guidance outperforms detailed procedural scaffolding, and that explicit task management tools have mixed effects on focused modeling tasks.
+The translation of natural language to formal constraint models requires expertise in the problem domain and modeling frameworks.
+To explore the effectiveness of agentic workflows, we propose CP-Agent, a Python coding agent that uses the ReAct framework with a persistent IPython kernel.
+We provide the relevant domain knowledge as a project prompt of under 50 lines.
+The algorithm works by iteratively executing code, observing the solver’s feedback, and refining constraint models based on execution results.
+We evaluate CP-Agent on 101 constraint programming problems from CP-Bench.
+We made minor changes to the benchmark to address systematic ambiguities in the problem specifications and errors in the ground-truth models.
+On the clarified benchmark, CP-Agent achieves perfect accuracy on all 101 problems.
+Our experiments show that minimal guidance outperforms detailed procedural scaffolding.
+Our experiments also show that explicit task management tools can have both positive and negative effects on focused modeling tasks.
 1 Introduction
-Constraint Programming (CP) is a declarative paradigm for solving combinatorial problems where users specify what constraints a solution must satisfy, rather than prescribing how to find it [26].
-The primary limitation preventing wider adoption of CP remains the modeling bottleneck: translating natural language problem descriptions into formal constraint models requires deep expertise in both the problem domain and the modeling framework [5].
-This challenge has been characterized as the “holy grail” of constraint programming [24].
-Large Language Models (LLMs) are now being applied to automate this translation process.
-The CP-Bench benchmark [14], introduced to evaluate LLM-driven constraint modeling capabilities, contains 101 diverse combinatorial problems sourced from various established CP community resources.
-These problems encompass optimization and satisfaction tasks, representing a variety of distinct constraint types.
-This benchmark addresses limitations of existing datasets, such as NL4Opt [17] and Logic Grid Puzzles [25], which lack the diversity and complexity necessary for evaluating constraint modeling capabilities.
-The CP-Bench study [14] evaluated multiple LLMs across three modeling frameworks.
-With direct prompting, the best models achieved up to 65% accuracy using Python-based frameworks (CPMpy, OR-Tools), while MiniZinc (a popular constraint modeling language) peaked around 50%.
-Applying inference-time methods (repeated sampling combined with self-verification) to CPMpy improved accuracy to approximately 70%.
-The choice between fixed and agentic workflows is well-established in the literature.
-As Wei et al.
-[29] demonstrated with chain-of-thought prompting, and subsequent work by Yao et al.
-[32] and Schick et al.
-[19] confirmed that one should use a fixed, single-shot workflow for well-scoped, short-horizon tasks, and prefer an agentic, tool-using loop as task complexity, multi-step reasoning, or external information needs increase.
-We hypothesize that constraint modeling from natural language falls into the latter category, requiring an agentic workflow with iterative refinement, hypothesis testing, and dynamic adaptation based on solver feedback.
-Our results confirm this hypothesis.
-To test this hypothesis, we introduce CP-Agent, a constraint modeling agent based on our general-purpose agentic-python-coder system.
-CP-Agent operates within a persistent IPython kernel [16], enabling stateful and iterative development.
-Rather than embedding domain-specific logic into its architecture, the agent is guided by a concise project prompt.
-This design allows the agent to execute code, observe runtime feedback from the constraint solver, and adjust models based on execution results.
-The underlying coding agent is also available as an MCP server [1].
-The Model Context Protocol (MCP), now governed by the Linux Foundation, has become a de facto standard for connecting language models with external tools [8].
-This integration enables interactive constraint modeling within MCP-compatible clients such as Claude Desktop or Cursor, where users can guide the agent through complex modeling decisions.
+Constraint Programming is a declarative paradigm for solving combinatorial problems in which users specify the constraints a solution must satisfy, rather than prescribing how to find it [26].
+The main obstacle preventing wider adoption of CP remains the modeling bottleneck.
+Converting natural language problem descriptions into formal constraint models requires deep expertise in both the problem domain and the modeling framework [5].
+This challenge is referred to as the “holy grail” of constraint programming [24].
+In recent years, Large Language Models (LLMs) have been successfully applied to automate this translation process.
+We considered CP-Bench [14], a benchmark introduced to evaluate LLM-driven constraint modeling capabilities, which contains 101 diverse combinatorial problems sourced from various established CP community resources.
+These problems are used as benchmarks for optimization and satisfaction tasks, and include a variety of distinct constraint types.
+This benchmark addresses the limitations of existing datasets, such as NL4Opt [17] and Logic Grid Puzzles [25], which lack the diversity and complexity needed to evaluate constraint modeling capabilities.
+Michailidis et al.
+[14] evaluated several LLM-based workflows on CP-Bench.
+With direct prompting, Python-based frameworks (CPMpy, OR-Tools) achieve overall good accuracy up to 65%.
+Indeed, MiniZinc (a popular constraint modeling language) peaked at around 50%.
+Repeated sampling can be combined with self-verification to improve accuracy for Python-based frameworks to approximately 70%.
+The relationship between fixed and agentic workflows is now well understood.
+Chain-of-thought prompting [29] and subsequent work [32, 19] have been widely used in the literature.
+One distinguishes between fixed workflows that use well-scoped, short-horizon tasks and agentic workflows where the LLM runs in a loop and uses tools.
+We hypothesize that constraint modeling from natural language falls into the category of tasks that require an agentic workflow with iterative refinement, hypothesis testing, and dynamic adaptation based on solver feedback.
+Our experiments show that this is indeed the case.
+To test this hypothesis, we introduce CP-Agent, a constraint modeling agent that utilizes our general-purpose agentic-python-coder system.
+CP-Agent maintains a persistent IPython kernel [16], enabling stateful and iterative development.
+Instead of domain-specific logic embedded in its architecture, we use a concise project prompt as the guide.
+The agent can execute code, receive runtime feedback from the constraint solver, and adjust its constraint model accordingly.
+The underlying coding agent is available as an MCP server [1].
+The Model Context Protocol (MCP), now governed by the Linux Foundation, is widely used for connecting language models with external tools [8].
+With this integration, the CP agent can be used from any MCP-compatible client, including coding assistants like Cursor or AI chatbots like Claude Desktop.
 On the CP-Bench dataset of 101 problems (with added clarifications for some ambiguous problems and output format requirements), CP-Agent solves all problems correctly.
-While our primary reference language model is Claude Sonnet 4.5, we also tested this approach with several other models from four different providers.
-Swapping only the project prompt enables the agent to adapt to new domains, as confirmed by follow-up work [23] on logic programming modeling tasks.
-We provide an experimental analysis of an explicit task management tool within our agentic coding workflow.
-The results are mixed: while task tracking helped on some previously failing problems, it also broke problems that worked without it.
-The net effect was negative (6 fewer problems solved) with significantly increased token usage, suggesting that for focused, single-file generation tasks, the overhead can distract as often as it organizes.
+We use Claude Sonnet 4.5 as the primary reference language model, though we also tested this with other models from different providers.
+Just by changing the prompt while keeping the agent architecture, we can adapt the system to different domains.
+This is exemplified by follow-up work [23] on modeling tasks for logic programming.
+In addition, we provide an empirical evaluation of using a task management tool within our agentic coding workflow.
+The results are mixed: task management is helpful for some but counterproductive for others; the net effect was adverse, with significantly increased token usage, suggesting that the overhead can distract as often as it organizes.
 2 Related Work
-CP-Agent aligns with the emerging neurosymbolic paradigm where neural language understanding is combined with symbolic constraint satisfaction [18].
-The LLM-based agent translates natural language into formal constraints, while the CP solver provides a sound and complete search over the solution space.
-Recent work has explored LLM agents for software engineering tasks.
-SWE-agent [31] resolves GitHub issues through a file-system-based Agent-Computer Interface with commands for navigating and editing repository files.
-CP-Agent focuses instead on single-script generation using a stateful IPython kernel, eliminating file I/O overhead for this class of problems.
-CodeAct [28] argues for using executable Python code as a unified action space for LLM agents, showing that this approach outperforms more restrictive formats.
-Our work applies this principle to constraint modeling, demonstrating that a general action space can solve specialized tasks by utilizing focused prompts.
-The iterative refinement in CP-Agent relates to automated program repair work.
-ChatRepair [30] introduced conversational repair, where an LLM iteratively refines patches based on test failures.
-RepairAgent [3] treats repair as an autonomous task with tools for code search, patching, and testing.
-Our agent performs similar iterative refinement but for generative tasks, translating natural language to constraint programs, rather than repairing existing code.
-Benchmark quality affects evaluation validity.
-EvalPlus [13] demonstrated that popular benchmarks contain insufficient tests, with augmented test suites revealing previously undetected errors.
-ClassEval [4] addressed structural complexity through manual benchmark curation.
-Our work on CP-Bench addresses problem specification ambiguity, enabling more rigorous measurement of constraint modeling abilities.
+Our approach follows the neurosymbolic paradigm, combining neural AI with symbolic constraint satisfaction [18].
+The LLM-based agent converts natural language into formal constraints.
+The CP solver guarantees formal correctness of the computation.
+LLM agents have been explored for software engineering tasks.
+SWE-agent [31] introduced the Agent-Computer Interface, a file-system-based approach for resolving GitHub issues that allowed navigating and editing repository files.
+Our approach is based on single-script generation using a stateful IPython kernel, which eliminates file I/O overhead for this class of problems.
+CodeAct [28] promotes executable Python code as a unified action space, and finds that this approach still outperforms more restrictive formats.
+We apply this general approach to constraint modeling, where a general action space can solve specialized tasks using focused prompts.
+The iterative refinement that we use in CP-Agent is related to automated program repair.
+ChatRepair [30] proposed conversational repair, where one iteratively refines patches based on test failures.
+RepairAgent [3] considers repair as an autonomous task with tools for code search, patching, and testing.
+Our approach is similar in spirit to iterative code repair, where refinement is used to correct errors.
+Still, we apply it to generative tasks, translating natural language into constraint programs rather than repairing existing code.
+The validity of an evaluation depends on benchmark quality.
+EvalPlus [13] showed that popular benchmarks have insufficient tests, while augmented test suites reveal previously undetected errors.
+Only recently, ClassEval [4] proposed a new benchmark for structural complexity through manual curation.
+We address problem ambiguity in the original CP-Bench problems and propose a fixed version of the benchmark that allows more rigorous testing of constraint modeling capabilities.
 3 Agentic Python Coder
-CP-Agent is based on a generic Python coding agent equipped with a special prompt for constraint modeling.
-The coding agent implements a minimal yet capable coding agent using the ReAct framework [32].
-The system architecture separates domain-agnostic agent capabilities from task-specific knowledge, enabling rapid adaptation to new problem domains through prompt engineering rather than code modification.
-The implementation of the generic coding agent is available from a GitHub repository and via PiPy for installation via “pip install agentic-python-coder” [20].
+Our CP-Agent is based on a generic Python coding agent, equipped with a special prompt for constraint modeling.
+Our coding agent implements the basic idea of ReAct [32] (Reason and Act) in a way that is both minimal and capable.
+The architecture can be adapted to new problem domains through prompt engineering rather than code modification.
+By this separation, domain-agnostic agent capabilities remain distinct from task-specific knowledge.
+Our implementation is publicly available on GitHub and via PyPI for installation via “pip install agentic-python-coder” [20].
 3.1 Architecture
 Virtual execution space
-Unlike software engineering agents that navigate complex file trees, this architecture targets single-script development, where all state is contained in a single file.
-The agent operates in a virtual execution space provided by the IPython kernel [16], which maintains a persistent state across multiple code executions.
-Variables, functions, and imports remain accessible throughout a session, eliminating file I/O overhead while enabling rapid hypothesis testing.
-The agent operates interactively via a Read–Eval–Print-Loop (REPL), and benefits from immediate feedback to all code manipulations.
-This design is well-suited for constraint programming models and similar focused generation tasks.
-To manage this persistent state and guide the execution flow, we employ LangGraph [12].
-This framework manages the agent’s workflow, enabling the development of stateful agents through durable execution and memory management.
-The framework supports the ReAct pattern [32] of interleaved reasoning and acting, allowing the agent to interact with the execution environment and adjust solutions through iterative cycles.
-The agent consists of three main components: the ReAct loop controller, which operates the reasoning-action cycle; a minimal tool suite focused on code generation and execution; and a project prompt that provides domain-specific guidance.
-The ReAct loop (Figure 1) manages the iterative cycle of reasoning about the task, selecting and executing tools, and observing results until the agent produces a final solution.
+Unlike software engineering agents, which navigate complex file trees, we do not maintain multiple files; instead, development is focused on a single file.
+We process development, refinement, and execution in a virtual space provided by the IPython kernel [16], which runs as a subprocess.
+The kernel provides a persistent state across multiple code executions.
+This guarantees that variables, functions, and imports remain accessible throughout a session and enables rapid hypothesis testing without file I/O overhead.
+The agent proceeds in rounds, and each round consists of code manipulation followed by immediate feedback via the Read–Eval–Print Loop (REPL).
+This design is well-suited for developing constraint programming models and similar focused generation tasks.
+The ReAct agent is implemented using LangGraph modules.
+LangGraph provides an important link between workflow management and durable execution, enabling the development of stateful agents.
+This even prevails when we consider the ReAct pattern [32] as interleaved reasoning and acting, which allows the agent to interact with the execution environment and adjust solutions through iterative cycles.
+We will now explain the details of each component: the ReAct loop controller, which implements the reasoning-action cycle (using LangGraph [12]); three tools the agent can call; and a project prompt that provides domain-specific guidance.
+In each iteration of the ReAct loop, the agent reasons about the task, selects and executes tools, and observes results until a final solution is obtained and saved to a file (Figure 1).
 TaskReasonAct (Tools)Solutionexecutecompleteobserve
 Figure 1: The ReAct (Reason and Act) framework employed by the agent.
-Tool suite
-The agent operates with a minimal, focused tool set.
-The core tool is python_exec(code: str), which executes Python code in the persistent kernel environment and returns stdout/stderr for solver feedback.
-Each execution builds on the previous state, allowing the agent to develop solutions incrementally.
-Calling save_code signals completion and writes the final solution to disk; a maximum iteration limit prevents runaway execution.
-Kernel implementation
-The system launches an IPython kernel as a separate subprocess, providing process isolation while enabling state persistence.
-All Python state resides in this long-running kernel, creating a stateful coding environment similar to an interactive session where each execution builds upon previous ones.
+Tools
+The main tool is python_exec(code: str), which runs Python code in the persistent kernel environment and returns stdout/stderr for solver feedback.
+Each call to this tool builds incrementally upon the previous calls.
+The function save_code transforms the solution into a file on disk, signaling completion.
+One could omit the save_code tool since file access is also possible via python_exec(code: str).
+However, having a dedicated tool for code saving has the advantage of easier file-path handling and makes it clear to the agent that calling this tool marks the end of its session.
+Optionally, we give the agent access to a task management tool to keep track of tasks in the form of a “to-do list.” The tool call has the format todo_write(todos: List[Dict[str, Any]]), where todos is the task list, and each entry specifies task attributes.
+By construction, each task must contain id, content, status, and priority fields, and only one task can be marked as “in_progress” at any time.
 Dynamic dependency injection
-Rather than having the agent execute pip install statements for adding required Python libraries or relying on pre-built container images, the system implements a custom kernel manager that wraps the standard IPython kernel launch.
-Using the uv package manager [2], the system injects requested packages directly into an ephemeral environment at kernel startup.
-This approach provides process-level isolation: each session runs in a clean environment with only the specified dependencies, while the kernel maintains state persistence across executions.
-The design decouples the agent’s reasoning from dependency mechanics, thereby eliminating the latency associated with runtime package installation.
-Task management
-The agent can use a task management tool to create and maintain an in-memory list of tasks, commonly referred to as a “todo list.” The tool call uses the format todo_write(todos: List[Dict[str, Any]]).
-Each task must contain id, content, status, and priority fields, with only one task marked as “in_progress” at any time.
+We implement a custom kernel manager, which wraps the standard IPython kernel launch.
+Instead of having the agent run pip install commands or rely on prebuilt container images, we inject packages at startup.
+Using the uv package manager [2], we split package installation into an ephemeral environment that runs at kernel startup.
+The kernel resolves dependencies based on session configuration; each session runs in a clean environment with only the specified packages.
 LLM integration
-The agent connects to language models through OpenRouter, [^1] https://openrouter.ai providing access without direct API dependencies.
-This approach allows switching between different models by changing a single configuration parameter.
-The system streams responses with usage statistics and detailed tracking of tool calls.
-These logs provide insights into the agent’s problem-solving process and resource consumption.
+The LLM that drives the agent is provided via the OpenRouter API.
+This allows us to use models from different model-makers through a uniform API and configuration, and with a single API key.
+We record all the agent’s steps and tool calls in log files.
+These logs provide valuable insight into the agent’s problem-solving process and resource consumption.
 Prompt architecture
-The agent operates with a three-layer prompt hierarchy.
-The system prompt (approximately 100 lines) defines general coding agent behavior and tool usage patterns without task-specific specialization, instructing the agent on how to reason about code, execute iterative development cycles, and utilize available tools effectively.
-The project prompt provides domain-specific knowledge (such as constraint modeling), examples, and best practices relevant to a particular problem class.
-The task prompt contains the specific problem to solve, typically read from a task.md file in the working directory or provided directly via a command line parameter.
-This separation enables the same agent infrastructure to tackle diverse problem domains by swapping project prompts while maintaining consistent operational behavior.
+The agent operates with the following three-fold prompt hierarchy: the system prompt (approximately 100 lines) specifies how the agent can be used to reason about code and execute iterative development cycles, defining general coding behavior and tool usage patterns without task-specific specialization.
+The project prompt combines domain-specific knowledge (such as constraint modeling) with examples and best practices relevant to this particular problem class.
+The concrete task to solve is in the task prompt, typically read from a task.md file in the working directory or provided directly via a command line parameter.
+The same agent infrastructure can tackle diverse problem domains by simply replacing the project prompt.
 4 CP-Agent
-For constraint programming, the generic coding agent is equipped with a domain-specific project prompt and uses CPMpy (see below) as the modeling framework.
-The approach relies on the agent’s ability to use general coding capabilities guided by domain expertise expressed in the project prompt, rather than specialized constraint programming logic embedded in the architecture.
-The agent can test simplified problem instances, verify constraint satisfaction on example data, implement custom validation logic, and refine models based on solver feedback.
-This flexibility, combined with the persistent execution environment, enables the agent to discover and correct modeling errors that would escape detection in single-pass generation approaches.
+For constraint programming, the generic coding agent is equipped with a domain-specific project prompt that instructs it to use CPMpy as the modeling framework.
+No specialized constraint programming logic is embedded in the architecture.
+The agent can create and test simplified problem instances and verify the constraint model on example data.
+The agent can even implement custom validation logic by refining models based on solver feedback.
+This flexible workflow, together with the persistent execution environment, provides an excellent environment for discovering and correcting modeling errors that would escape detection in single-pass generation approaches.
 4.1 CPMpy
-The agent uses CPMpy [7] as the constraint programming foundation.
-This Python-based library builds on NumPy for efficient handling of decision variables as arrays, allowing integration with the broader Python ecosystem.
-Its solver-agnostic design enables modeling problems once and solving them using various backends, including CP, MIP, SMT, and SAT solvers.
-The choice of Python matches findings from Michailidis et al.
-[14] that LLMs perform better with Python-based frameworks than domain-specific languages, likely due to the billions of lines of Python code in their training data.
+We use CPMpy [7] as the constraint programming foundation primarily due to its Python integration.
+This library is based on NumPy for efficient handling of decision variables as arrays, making it well-suited for integration with the broader Python ecosystem.
+Its solver-agnostic design allows a more refined selection of backends and hence enables modeling problems once and solving them using various solvers, including CP, MIP, SMT, and SAT.
+This design is in line with findings from Michailidis et al.
+[14] that LLMs perform better with Python-based frameworks than with domain-specific languages, likely due to the billions of lines of Python code in their training data.
 4.2 Project Prompt
-The project prompt (cpmpy.md) is under 50 lines and provides focused guidance for constraint modeling. The complete prompt is included in the GitHub repository [20].
-The prompt contains five core rules: use CPMpy’s declarative modeling rather than implementing search algorithms, output only valid JSON using json.dumps(), always import the json module when outputting JSON, check the exact output format required, and manually test solutions to ensure they satisfy the problem requirements.
+The project prompt (cpmpy.md), under 50 lines, provides succinct guidance for constraint modeling.
+The complete prompt is available online at the GitHub repository [20].
+The prompt contains the following core rules: use CPMpy’s declarative modeling rather than implementing search algorithms, output only valid JSON using json.dumps(), always import the json Python module when outputting JSON, check the exact required output format, and manually test solutions to ensure they satisfy the problem requirements.
 A basic code template demonstrates the standard structure, and brief sections cover three essential CPMpy constraints (AllDifferent, sum, Circuit) and optimization via model.minimize() and model.maximize().
-This minimal prompt resulted from an ablation study.
-Initial experiments utilized a substantially longer prompt, approximately 800 lines in length, which contained detailed modeling procedures, extensive problem-type catalogs, and debugging guidance.
-The shorter version achieved equal or better performance, suggesting that for modern LLMs with constraint programming knowledge in their training data, concise guidance focusing on common pitfalls suffices—excessive detail may limit problem-solving flexibility.
+This minimal prompt is based on an ablation study.
+We initially experimented with a substantially longer prompt of approximately 800 lines, containing detailed modeling procedures, extensive problem-type catalogs, and debugging guidance.
+The shorter version performs on par with, or better than, the longer version across all performance metrics.
+This indicates that for modern LLMs with constraint-programming knowledge in their training data, concise guidance is sufficient; excessive detail may distract and consume input tokens without providing benefit.
 5 Benchmark Problems
-The CP-Bench dataset [14] provides a benchmark for evaluating automated constraint modeling capabilities.
-It contains 101 constraint programming problems drawn from four established sources: CSPLib [6], CPMpy examples [7], models by Kjellerstrand [11], and problems from the APLAI course at KU Leuven.
-The problems span diverse combinatorial classes, including scheduling, assignment, sequencing, graph problems, and combinatorial design, with 30 optimization problems and 71 decision problems providing balanced coverage of constraint programming challenges.
-Each problem is presented as a natural language description in Markdown format, along with input data and required output variables.
-Solutions must execute within a specified time limit and produce results in JSON format, using the specified variable names.
-CP-Bench includes ground-truth CPMpy models for all problems, enabling semantic validation through constraint satisfaction checking rather than relying solely on syntactic correctness or execution success.
-This semantic validation is essential for rigorous evaluation of automated modeling approaches.
-The ground-truth models are withheld from the agent and used solely for post-hoc verification of generated solutions.
+The CP-Bench collection [14] provides an excellent benchmark for testing automated constraint modeling capabilities.
+This collection comprises 101 constraint programming problems from four established sources: CSPLib [6], CPMpy examples [7], models by Kjellerstrand [11], and problems from the APLAI course at KU Leuven.
+The collection covers diverse combinatorial problem types, including scheduling, assignment, sequencing, graph problems, and combinatorial design.
+With 30 optimization and 71 decision problems, the collection provides a balanced coverage of constraint programming challenges.
+Each problem in CP-Bench consists of a natural-language description in Markdown format, input data, required output variables, and a ground-truth model for verification.
+Solutions must output results in JSON format using the specified variable names.
+The ground-truth models enable semantic validation through constraint satisfaction checking rather than syntactic correctness or execution success.
+The ground-truth models are withheld from the agent and used solely for post-hoc validation.
 5.1 Fixed Workflow Results
 Michailidis et al.
-[14] evaluated fixed workflow approaches on CP-Bench, testing progressively detailed prompts (basic instructions, modeling guidelines, API documentation) and retrieval-augmented in-context learning.
-The highest accuracy of 70% was achieved by combining repeated sampling with self-verification using GPT-4-turbo.
-These fixed workflow approaches share a common limitation: they cannot test hypotheses, debug errors, or verify partial solutions.
+[14] evaluated fixed workflow approaches on CP-Bench, testing progressively detailed prompts including basic instructions, modeling guidelines, and API documentation, as well as retrieval-augmented in-context learning.
+Their results indicate that combining repeated sampling with self-verification using GPT-4-turbo achieves the highest accuracy of 70%.
+A general limit of these fixed workflow approaches is that they do not support testing hypotheses, debugging errors, or verifying partial solutions.
 5.2 Benchmark Corrections and Clarifications
-Our initial evaluation attempts revealed issues with the original benchmark that prevented rigorous automated evaluation.
-The CP-Bench authors independently acknowledged222personal communication, August 2025 several problems, including ground-truth models with symmetry-breaking constraints that reject valid solutions, unclear output format requirements, and ambiguities in problem statements.
-Without clear specifications, evaluation conflates the ability to model constraint satisfaction problems with the ability to guess intended formats and interpret ambiguous descriptions.
-We added clarifications to isolate the actual modeling capability.
-These clarifications are necessary to maintain focus on constraint modeling rather than resolving ambiguity or making guesses.
+During testing, we encountered issues with some CP-Bench problems.
+The CP-Bench authors acknowledged these issues independently111personal communication, August 2025, including ground-truth models with symmetry-breaking constraints that reject valid solutions, unclear output format requirements, and ambiguities in problem statements.
+Without a clear problem specification, the modeling ability cannot be evaluated separately from format guessing.
+We therefore updated the CP-Bench problems to address these issues.
 Our benchmark updates fall into three categories.
 First, we clarified 31 ambiguous problem statements.
-For instance, Problem 080 (TSP) required specifying whether Euclidean distances should be rounded or truncated—different methods lead to different distance matrices and thus different optimal solutions.
+For instance, Problem 080 (TSP) required specifying whether Euclidean distances should be rounded or truncated, which can lead to different distance matrices and thus different optimal solutions.
 For Problem 013 (magic square), we corrected a misleading description that stated the goal is to find a “magic square” when the problem actually requires a “magic sequence.” Problem 060 (crossword) provided a visual grid and word list, but required explicit specification of slot lengths and intersection constraints to be solvable.
 Second, we added structured JSON schemas to all problems to clarify the required output format.
-Third, we corrected errors in 19 ground-truth models: removing symmetry-breaking constraints from 10 and fixing logic errors in 9 others.
-The modifications do not make the constraint modeling problems easier—they remove confounding factors that obscure actual modeling capability.
-The underlying combinatorial structure and search space complexity remained unchanged; no changes hint at constraint formulations or modeling strategies.
-The success rates reported in the original CP-Bench paper were obtained using the original benchmark and ground-truth models.
-Direct comparison between our results and theirs would be inappropriate, as they measure different capabilities.
-With clear specifications, other methods would likely show higher success rates as well.
-We document all changes in supplementary materials and provide the clarified benchmark for fair comparison in future work.
-6 Experiment
+We recommend this practice for any future extensions of the benchmark collection, as it eliminates trivial but annoying discrepancies, such as whether array indices are supposed to start with 0 or 1.
+Third, we corrected errors in 19 ground-truth models: we removed symmetry-breaking constraints from 10 and fixed logic errors in 9 others.
+While these updates do not make the problems easier, they remove confounding factors that obscure actual modeling capability.
+The underlying combinatorial structure and search space complexity remain the same; no changes hint at constraint formulations or modeling strategies.
+The success rates reported in the original CP-Bench paper were obtained without using our clarifications.
+A direct comparison between our results and theirs is not possible, as the fixed workflow can also benefit from the clarifications.
+The clarified benchmark is available online [22].
+6 Experiments
 6.1 Experimental Setup
-Our experimental setup executed all 101 problems using CP-Agent, implemented as agentic-python-coder (version 2.2.1), with Claude Sonnet 4.5 (version 20250929).
-Each problem execution produces a solution file containing the generated CPMpy model and a conversation log recording the complete agent interaction.
-We validated the solutions by running them against the ground-truth CPMpy models.
-For satisfaction problems, we verified that the solutions satisfy all ground-truth constraints.
-For optimization problems, we also compared the objective value to confirm optimality.
-This approach correctly handles problems with multiple valid solutions while ensuring correctness and accuracy.
-We ran each experiment three times to account for stochastic variance.
-The agentic-python-coder and all experimental data, including problem specifications, ground-truth models, generated solutions, and execution logs, are available online [22].
-The data is encrypted (password Ckc1ADsKXRfmzD2) to prevent training data contamination.
+We ran all our experiments on all 101 CP-Bench problems three times, with CP-Agent implemented as agentic-python-coder (version 2.2.1) using Claude Sonnet 4.5 (version 20250929) as the language model for our main experiment.
+The models used in the model generalization and the task management experiments are described below.
+For each problem, the agent produced a solution file containing the generated CPMpy model, and we recorded a conversation log with the complete agent interaction.
+We tested all solutions against the ground-truth CPMpy models.
+For each satisfaction problem, we tested whether the solution satisfies all ground-truth constraints.
+For optimization problems, we compared the objective value of the generated solution with an optimal pre-computed ground-truth value to confirm optimality.
+Note that a problem may have several correct and optimal solutions, hence the ground-truth model must not contain symmetry-breaking constraints.
+We always report results averaged across three runs to account for randomness.
+The agentic-python-coder, the problem specifications and ground-truth models to reproduce our experiments, and the generated solutions and execution logs, are available online [22].
+The data is encrypted (password Ckc1ADsKXRfmzD2) to avoid training data contamination.
 6.2 Results
-CP-Agent achieved 100% accuracy on all 101 CP-Bench problems, successfully solving all problems across all three independent runs.
-All generated solutions were validated against the reference CPMpy models and produced correct and optimal solutions.
-Table 1: Tool usage, token consumption, and execution time for all 101 CP-Bench problems (averaged over three independent runs).
+Our experimental results show that CP-Agent achieves 100% accuracy, successfully solving all 101 CP-Bench problems across all three independent runs.
+We validated all generated solutions against the ground-truth models and confirmed that all solutions were correct and optimal.
+Our results are summarized in the table on the next page: tool usage, token consumption, and execution time for all 101 CP-Bench problems (averaged over three independent runs).
 Column ex shows the number of python_exec calls (iterative refinement cycles).
@@
 Column time shows execution time in seconds (rounded).
+The number of python_exec calls varies and depends on the structure of the problem, ranging from 1 to 28, with simpler problems requiring fewer iterations and complex scheduling or optimization problems requiring more exploration.
+Problem 014 (crossfig) stands out with an unusually large number of Python executions (28 average), followed by 009 (basketball) with 24 executions, reflecting their combinatorial complexity.
 Problem ex in out time Problem ex in out time
@@
 051_circling_s 2.3 25 3 47
-Token consumption averaged 37k input and 3k output tokens per problem.
-Problem 014 (crossfig) represents an extreme outlier, consuming 737k input tokens due to its complex crossword-figure structure, which requires extensive iterative constraint refinement.
-Execution times ranged from 20 seconds to 542 seconds, with an average of 52 seconds per problem across the three runs.
-The timing variation between the three independent runs (averaging 48, 57, and 52 seconds, respectively) demonstrates the stochastic nature of LLM-based systems.
-Table 1 summarizes computational cost; solution quality is perfect across all runs.
-The agent called python_exec between 1 and 28 times per problem, with simpler problems requiring fewer iterations and complex scheduling or optimization problems requiring more exploration.
-Problem 014 (crossfig) required the most Python executions (28 average), followed by 009 (basketball) with 24 executions, reflecting their combinatorial complexity.
-To illustrate how the agent approaches constraint modeling, we examine two representative runs that demonstrate different problem-solving patterns.
+The table also shows the token consumption.
+On average we have 37k input and 3k output tokens per problem.
+As expected, Problem 014 (crossfig), the problem with the largest number of exec calls has also the largest token consumption (737k input tokens).
+The specifics of execution times differ from problem to problem, ranging from 20 seconds to 542 seconds, with an average of 52 seconds across the three runs.
+The average run time per problem in each of the three runs is 48, 57, and 52 seconds, respectively, which demonstrates the stochastic nature of LLM-based systems, but can also reflect variety in API response time and network speed.
+We present representative runs on two problems to illustrate how the agent approaches constraint modeling, with different problem-solving patterns.
 6.3 Problem 061: Cryptarithmetic Puzzle
-The cryptarithmetic problem required assigning unique digits (0–9) to letters in a 20-digit addition equation, ensuring the arithmetic holds while each letter maps to a distinct digit.
-The agent’s solution process demonstrates strategic adaptation when an initial approach proves unworkable.
-The first four executions built the foundation: defining ten integer variables (one per letter), adding an AllDifferent constraint, mapping variables to their positions in the three numbers, and constraining leading digits to be non-zero.
-This setup proceeded smoothly, establishing the basic problem structure.
-The fifth execution attempted to model the problem as a single unit by treating each 20-digit number as a single large integer variable.
-The agent constructed expressions like num1 = letter[0]*10**19 + letter[1]*10**18 + ...
-for all three numbers and added a constraint num1 + num2 == result.
-When the model tried to solve, the solver raised a TypeError: __init__(): incompatible constructor arguments because variable domains with bounds near $10^{20}$ exceeded the solver’s integer representation capacity.
-Recognizing the fundamental incompatibility between this modeling approach and solver limitations, the agent abandoned the large-integer strategy entirely.
-The agent redesigned the model, introducing binary carry variables and modeling addition column by column, thereby mimicking manual arithmetic:
+The problem is to assign letters with unique digits (0–9) such that the given 20-digit addition equation holds. The log reveals how the agent adapts when an initial approach proves unworkable.
+In the first four executions, the agent introduces integer variables (one per letter), an AllDifferent constraint, mappings from variables to their positions in the three numbers, and constraints forcing leading digits to be non-zero.
+This completes the description of the problem structure.
+The fifth execution attempts to treat each 20-digit number as a single large integer variable.
+The agent constructs expressions like num1 = letter[0]*10**19 + letter[1]*10**18 + ...
+for all three numbers and adds the constraint num1 + num2 == result.
+This approach fails because variable domains of size $10^{20}$ exceed the solver’s integer representation capacity, causing a TypeError: __init__(): incompatible constructor arguments.
+With this feedback, the agent redesigns the model, introducing binary carry variables and models addition column by column, thereby mimicking manual arithmetic:
 carries = intvar(0, 1, shape=21)
@@
 == result_digit + 10 * carry_out)
-This column-wise formulation ensured that all intermediate values remained within standard integer bounds, enabling the solver to determine the correct digit assignment.
+This way, all intermediate values are kept within standard integer bounds, enabling the solver to determine the correct digit assignment.
 6.4 Problem 082: Wolf, Goat, and Cabbage
-While the above example demonstrates error recovery, this river-crossing puzzle illustrates the agent’s behavior when proceeding without failures.
-The agent built the temporal model through four distinct phases.
-The first phase established state variables: wolf_pos[t], goat_pos[t], cabbage_pos[t], and boat_pos[t], which track locations (shore 0 or 1) across an 8-step timeline.
-The second phase fixed boundary conditions—all entities start at shore 0 and end at shore 1.
-The third phase encoded the movement mechanics: the boat crosses at each step, the farmer transports at most one item per trip, and item movements coincide with boat movements.
-The fourth phase added safety constraints:
+The above example demonstrated error recovery; this well-known river-crossing puzzle illustrates the agent’s behavior when proceeding without failures.
+The first step is to establish state variables: wolf_pos[t], goat_pos[t], cabbage_pos[t], and boat_pos[t], which track locations (shore 0 or 1) across an 8-step timeline.
+The second step of the construction fixes boundary conditions—all entities start at shore 0 and end at shore 1.
+The third step encodes the movement mechanics: the boat crosses at each step, the farmer transports at most one item per trip, and item movements coincide with boat movements.
+The fourth step adds safety constraints:
 for i in range(steps):
@@
 & (wolf_pos[i] != boat_pos[i]))
-These constraints forbid the goat and cabbage (or wolf and goat) from sharing a shore without the farmer present.
-The systematic layering—state, boundaries, movement, and safety—reflects a deliberate construction process where the agent tests each phase before proceeding to the next.
-The model solved correctly on the first execution.
-These two problems reveal complementary behaviors: complete redesigns when facing fundamental obstacles (Cryptarithmetic) and methodical incremental construction when the path is clear (Wolf, Goat, and Cabbage).
-The agent adapts its approach rather than following a fixed procedure.
+These constraints ensure that the goat and cabbage (or wolf and goat) do not share a shore without the farmer present.
+The systematic layering—state, boundaries, movement, and safety—naturally reflects the construction process.
+Each phase is traversed in order to verify correctness before proceeding to the next.
+The execution completes successfully on the first attempt.
+These two behaviors of the agent can be observed in different phases of problem solving: complete redesigns when facing fundamental obstacles and methodical incremental construction when the path is clear.
+The agent uses adaptive approaches rather than fixed workflows.
 7 Model Generalization
-To evaluate whether the approach generalizes beyond Claude Sonnet 4.5, we tested six frontier models on ten representative problems with three runs each.
-The models span four providers: Anthropic (Sonnet 4.5, Opus 4.5, Haiku 4.5), OpenAI (GPT-5.1-codex-mini), xAI (Grok-code-fast-1), and DeepSeek (v3.2).
-We excluded Google’s Gemini models due to infrastructure issues during the experimental period.
-All models used a temperature of 0.0 with a 20-minute timeout per problem.
-Table 2 shows the results.
-All tested models achieved a success rate of 80–100% using this architecture.
-Sonnet 4.5 maintained perfect accuracy, while Opus 4.5 and GPT-5.1 each reached 97%.
-Haiku 4.5, Grok, and DeepSeek achieved 80–87% accuracy, with most failures occurring on Problem 009 (basketball), although sporadic failures also occurred on other problems.
+To study whether the approach generalizes beyond Claude Sonnet 4.5, we evaluated six frontier models on ten representative problems, with three runs per model.
+The models come from four model makers: Anthropic (Sonnet 4.5, Opus 4.5, Haiku 4.5), OpenAI (GPT-5.1-codex-mini), xAI (Grok-code-fast-1), and DeepSeek (v3.2).
+Due to infrastructure issues during the experimental period, we did not test with Google’s Gemini models.
+We ran each problem with a 20-minute timeout, using a temperature of 0.0 for each model.
+We give the results in Table 1.
+The tested models achieved a success rate of 80–100%.
+Sonnet 4.5 maintained perfect accuracy.
+Opus 4.5 and GPT-5.1 both reached 97%.
+Haiku 4.5, Grok, and DeepSeek achieved 80–87% accuracy; most failures occurred on Problem 009 (basketball), though sporadic failures also affected other problems.
 Basketball requires coordinating nine constraint categories, including mirroring schemes, home/away patterns, rival pairings, and date-specific requirements—the highest constraint density in this subset.
-Some runs hit the 20-minute timeout before completion.
-Review of execution logs confirms these agents were not converging toward correct solutions, suggesting the timeout did not cut short otherwise successful attempts.
-These results show the architecture generalizes across model families.
-Success correlates with model capability rather than provider: the agentic workflow provides value at all capability levels tested.
-Prompt optimization frameworks, such as DSPy [10], could automatically tune the project prompt for specific models, potentially narrowing the performance gap observed with other models.
+For some runs, the solver timed out after 20 minutes.
+Looking at execution logs, we see that these agents were not converging toward correct solutions, suggesting the timeout did not cut short otherwise successful attempts.
+Our results show that the architecture generalizes across model families.
+This indicates that our agentic workflow generalizes from our default model, Sonnet.
+Opus, considered superior to Sonnet, showed slightly weaker performance, possibly because the prompt was optimized for Sonnet.
+Prompt optimization frameworks such as DSPy [10] could be used to automatically tune the project prompt for specific models, potentially narrowing the performance gap observed with other models.
 Table 1: Multi-model evaluation (3 runs each). Numbers indicate successful runs out of 3. Model success rates: Sonnet 100%, Opus/GPT-5 97%, Haiku 87%, DeepSeek 83%, Grok 80%.
@@
 Explicit task tracking for LLM agents—maintaining a mutable list of pending, in-progress, and completed tasks—was introduced by BabyAGI [15] and formalized in plan-and-solve prompting [27].
-The todo_write tool in our coder implements this pattern.
+Our implementation provides a tool that realizes this pattern.
 Our architecture requires only todo_write, not a corresponding todo_read: without history compaction, all previous tool calls remain in the conversation context, so the agent can always see its earlier task updates.
-To evaluate whether explicit task management benefits constraint modeling, we conducted an ablation study comparing performance with and without this tool.
-We used a smaller model (Claude Haiku 4.5) than our default model (Claude Sonnet 4.5) to reveal potential differences.
-We selected 40 problems, stratified by difficulty: 20 easy problems (solved in 27 seconds or less by Sonnet 4.5) and 20 hard problems (requiring 58 seconds or more).
-For each problem, we executed three independent runs with todo_write disabled and three with it enabled.
-Table 3 presents the results.
+To evaluate whether explicit task management benefits constraint modeling, we compare our method with and without this tool.
+For this experiment, we use a smaller model (Claude Haiku 4.5) than our default (Claude Sonnet 4.5) to reveal potential differences.
+We consider the following 40 problems, stratified by difficulty: 20 easy problems (solved in 27 seconds or less by Sonnet 4.5) and 20 hard problems (requiring 58 seconds or more).
+We run each problem three times with todo_write disabled and three times with it enabled.
+The results are summarized in Table 2.
 Table 2: Task management ablation on 40 problems (Claude Haiku 4.5, 3 runs per condition).
@@
 Easy (60 runs) 59 60 Hard (60 runs) 52 45
-On easy problems, todo_write increased token consumption by 191% (from 15K to 44K average) and execution time by 111% (from 18s to 38s), while improving success by just one run (59/60 to 60/60).
-On hard problems, todo_write reduced success from 52/60 to 45/60 while still adding overhead: 41% more tokens (209K to 295K) and 27% longer execution (98s to 124s).
-Examining individual problems reveals the pattern.
-The task management tool helped three previously failing hard problems—crossfig, sports, and progressive—by providing structure for extended problem-solving.
-However, it broke three problems that worked without it: basketball, crew, and wolf_goat all dropped from 3/3 to 0/3.
-The task-tracking overhead appears to distract the agent from solutions it would otherwise find.
-9 Discussion
-Our experiments yield several key observations regarding the design and evaluation of agentic systems for constraint modeling.
-The ablation study, which shows that a short prompt outperforms an 800-line version, suggests that for foundation models with constraint programming knowledge in their training data, the prompt’s role is to set high-level strategy and details on output format rather than provide exhaustive procedural instructions.
-Prompt compression research supports this finding: shorter, denser prompts can match or exceed verbose prompts [9].
-Excessive detail may limit the agent’s ability to solve problems effectively.
-The primary evaluation used Claude Sonnet 4.5.
-Multi-model testing on a representative 10-problem subset confirms that the architecture generalizes: six models from four providers achieved success rates of 80–100%, with performance correlating with model capability.
-This pattern extends to built-in planning mechanisms.
-The optional todo_write tool, intended to help the agent manage tasks, reduced overall success when tested with a smaller model.
-The effect was asymmetric: todo_write helped on already-failing problems but broke problems that worked without it.
-For problems where the entire solution fits in a single script, the overhead of maintaining an explicit task list is more distracting than beneficial.
-When state is fully captured in code and execution output, the agent’s internal reasoning seems to suffice without externalized planning.
+On easy problems, compared with the baseline, todo_write results in a significant increase in token consumption (191%, from 15K to 44K average) and execution time (111%, from 18s to 38s), while improving success by just one run.
+On hard problems, todo_write even reduced success from 52/60 to 45/60 while still adding overhead in terms of token usage and longer execution time.
+A closer look at individual problems reveals a pattern.
+The task management tool provides structure for extended problem-solving, which helps three previously failing hard problems—crossfig, sports, and progressive.
+In contrast, it causes failures on three problems that worked without it.
+The task-tracking overhead appears to distract the agent from finding solutions in some cases.
 9 MCP Server
-The Model Context Protocol (MCP) provides a standardized interface for connecting language models with external tools and data sources [8].
-Introduced by Anthropic in late 2024 [1], MCP has become a de facto standard for agent-tool interaction, with adoption by major AI platforms including OpenAI and integration into development environments such as Cursor, Claude Code, VS Code (via Cline), Windsurf, and OpenAI Codex CLI.
-In December 2025, MCP governance transferred to the Agentic AI Foundation under the Linux Foundation.
-While our standalone CP-Agent excels at batch processing, interactive development benefits from human guidance through complex modeling decisions.
-We expose the same persistent IPython kernel as an MCP server (ipython_mcp) for integration with Model Context Protocol-compatible clients such as Claude Desktop or Cursor.
-The server provides python_exec for code execution, the same tool central to our standalone agent, along with three tools for controlling and monitoring the coding infrastructure: python_reset to clear state and install packages to the virtual environment, python_status to query session information, and python_interrupt to halt long-running computations.
+The Model Context Protocol (MCP) is a standardized interface for connecting language models with external tools and data sources [8].
+MCP was initially introduced by Anthropic in late 2024 [1] as a tool for agent-tool interaction, and has since been adopted by major AI platforms and integrated into development environments.
+In December 2025, MCP governance transferred to the Linux Foundation.
+To make our framework accessible from other agentic systems, beyond our own ReAct agent, we define the same persistent IPython kernel system as an MCP server (ipython_mcp).
+This allows it to be integrated with any MCP-compatible client.
+The MCP server is equipped with python_exec for code execution, the same tool central to our standalone agent.
+It also comprises three tools for controlling the coding infrastructure: python_reset to clear state and install packages, python_status to query session information, and python_interrupt to halt long-running computations.
 This interactive mode also serves an educational purpose: users can observe how natural language constraints are formalized into CPMpy models, request explanations, and refine models based on solver feedback.
-Once developed, the generated code can be saved for deployment in other contexts.
-Our MCP integration complements the system MCP-Solver [21], which provides specialized tools for constraint modeling (MiniZinc, PySAT, Z3, ASP).
-In contrast, the ipython_mcp server offers general-purpose Python execution, where domain adaptation occurs through prompts rather than tool specialization.
+Once developed, the generated code can be deployed to other contexts.
+This approach complements the MCP-Solver [21], which provides specialized tools for constraint modeling (MiniZinc, PySAT, Z3, ASP).
+In contrast to the MCP-Solver, our new ipython_mcp server provides general-purpose Python execution, with domain adaptation driven by prompts rather than tool specialization.
 10 Conclusion
-CP-Agent solves all 101 problems in the clarified CP-Bench benchmark, demonstrating that agentic workflows with execution feedback can reliably automate constraint modeling from natural language.
-The agent architecture combines the ReAct framework with a persistent IPython kernel.
-Our ablation study shows that a minimal project prompt is sufficient, and that explicit task management tools generally add overhead for focused modeling tasks.
-However, they can help maintain focus during longer problem-solving sessions.
-For foundation models with relevant domain knowledge, concise guidance and stateful execution prove more effective than detailed procedural scaffolding or explicit planning mechanisms.
-The architecture’s separation of domain expertise (prompts) from execution infrastructure (the agent) allows adaptation to other scientific computing domains.
-Follow-up work on ASP-Bench [23] demonstrated excellent performance on 128 answer set programming problems, utilizing the identical architecture with only a domain-specific project prompt, thereby confirming that our approach generalizes across declarative programming paradigms.
-Acknowledgments
-This research was supported by the Austrian Science Fund (FWF) projects 10.55776/P36420 and 10.55776/COE12.
+Our experiments provide several insights into the design and evaluation of agentic systems for constraint modeling.
+Our results show that agentic workflows with execution feedback can reliably automate constraint modeling from natural language.
+Our main experiment demonstrates that a short prompt performs very well (even outperforms an 800-line earlier version).
+This suggests that, for foundation models with constraint-programming knowledge, the role of the prompt is to set high-level strategy and the output format rather than to provide exhaustive procedural instructions.
+This is in line with recent results on prompt compression showing that shorter, denser prompts can match or exceed verbose prompts [9].
+In terms of procedural detail, excessive instruction might reduce the agent’s ability to solve problems effectively.
+Our model generalization experiments show that our agentic architecture is robust: six models from four providers achieved success rates of 80–100%, with performance correlating with model capability.
+Our experiment with the task management tool gives a mixed picture.
+While the tool reduces the overall success rate and significantly increases time and token usage, it was helpful for a few problems.
+Based on this outcome, we turn off the tool by default but keep it in the codebase for further testing, as there might be problem families where it shows a positive overall performance.
+Finally, we reported on further integration possibilities of our system via the Model Context Protocol.
+We believe that separating domain expertise from execution infrastructure, as realized in our system, can be beneficial in other scientific computing domains.
+In follow-up work, we used our system to transform natural language descriptions into logic programs [23], using the same architecture with only a domain-specific project prompt.